OpenStudy (anonymous):

Hi I need help with my home work can someone please I need to describe the whole number vaules of x that make the statement true. x divided by 2/3 =x

Directrix (directrix):

The set of whole numbers is the set comprised of {0, 1, 2, 3, 4, ...} x divided by 2/3 =x x/(2/3) = x Looking at this equation, think "what number is equal to 2/3 of itself?" If the answer does not come to you, then use Algebra. x/(2/3) = x x/(2/3) - x = 0 x [ (2/3) - 1] = 0 x (-1/3) = 0 x = 0 / (-1/3) x = 0 which is a whole number.